# Towards Scalable Multi-Chip Wireless Networks with Near-Field Time Reversal

### v2

@@ Abstract
The concept of Wireless Network-on-Chip (WNoC) has emerged as a potential solution to address the escalating communication demands of modern computing systems due to its low-latency, versatility, and reconfigurability. However, for WNoC to fulfill its potential, it is essential to establish multiple high-speed wireless links across chips. Unfortunately, the compact and enclosed nature of computing packages introduces significant challenges in the form of Co-Channel Interference and Inter-Symbol Interference, which not only hinder the deployment of multiple spatial channels but also severely restrict the symbol rate of each individual channel. In this paper, we posit that Time Reversal (TR) could be effective in addressing both impairments in this static scenario thanks to its spatiotemporal focusing capabilities even in the near field. Through comprehensive full-wave simulations and bit error rate analysis in multiple scenarios and at multiple frequency bands, we provide evidence that TR can increase the symbol rate by an order of magnitude, enabling the deployment of multiple concurrent links and achieving aggregate speeds exceeding 100 Gb/s. Finally, we evaluate the impact of reducing the sampling rate of the TR filter on the achievable speeds, paving the way to practical TR-based wireless communications at the chip scale.
Index Terms:
Time Reversal; Near-Field Communications; Wireless Interconnects; Network-on-ChipI Introduction
@@
In this context, Wireless Network-on-Chip (WNoC) is an emerging paradigm with significant promise to enhance communications within computing packages. WNoC are possible because of the advances in antennas and transceivers in the millimeter-wave (mmWave) and terahertz (THz) bands that have made possible their integration inside computing cores [3, 4]. They are expected to mitigate the scalability and flexibility issues that come with the increment of the number of processors, their integration on multiple interconnected chips, and the use of wired NoC communication approaches [1, 5]. The goal is to augment the wired backbone with wireless links within the package [6].
@@
Due to the use of electromagnetic propagation nearly at the speed of light, WNoC provides broadcast and multicast opportunities with reduced latency and is expected to offer reconfigurable links able to operate at speeds of tens of Gb/s or more. For the WNoC paradigm to be competitive, though, it must deliver a reliability commensurate to that of their wired counterpart. In particular, multiple works [7, 8, 9, 10] point towards a latency of 1–10 nanoseconds, a throughput in the range of 10–100 Gb/s or higher, and a Bit Error Rate (BER) of 10-12–10-15 as targets for WNoC. Complying with these values while keeping low the area/power consumption calls for affordable solutions, low order modulations, and simple physical layer elements for implementation [11, 12, 9, 13, 14, 15].
@@
However, several challenges arise when including wireless communications in a package environment while trying to meet the target performance requirements. Some examples include the antenna placement in the highly integrated environment, design of transceivers with efficient coding schemes and modulations, interference management, and medium access control [6, 16, 17, 18]. In particular, in this paper we focus on the problems of Inter-Symbol Interference (ISI) and Co-Channel Interference (CCI). In the on-chip scenario, the antennas radiate in a highly reverberant environment and are impaired by near-field effects. As a result, on the one hand, signals are subject to multiple reflections before reaching the receiver, leading to large delay spreads dominated by the strong multipath components. This leads to severe ISI, which degrades the performance of high-rate transmissions especially in low-order modulations. On the other hand, reverberation causes the radiated energy to spread across the entire space. This leads to strong CCI with non-intended antennas and prevents the possibility of having multiple spatial channels concurrently. Other techniques such as Orthogonal Frequency Division Multiple Access (OFDMA) and beamforming could address ISI and CCI, but are not advisable in the chip-scale communications scenario due to the required complex and power-hungry transceivers, as well as the need for multiple antennas consuming a large area, respectively [19, 20].
@@
Time Reversal (TR) [21, 22, 23, 24, 25] is a technique that uses a time-reversed version of the channel’s impulse response (CIR) to create a matched filter that focuses energy in space and time in a targeted destination. Due to the spatiotemporal focusing, TR can therefore considerably alleviate both ISI and CCI, as shown in multiple works at multiple scales and including in the near-field [26, 22].
@@
The rest of the paper is organized as follows. In Section II, we provide some background related to the general problems of the on-chip scenario. In Section III, we formulate the system model and in Section IV, we provide a detailed description of the methodology, including the simulation pipeline and the evaluated variations. We present the results obtained for the parametric characterization from the base scenario in Section V, while Section VI evaluates the performance of TR with the changes in the antenna operating frequency and the geometry of the chip. Finally, in Section VII, we assess the trade-off between performance and feasibility of TR filters.
@@ II-A Wireless Communications at the Chip Scale
Wireless on chip communications have been built with the foundation of using mmwave technology, as an overlay of wireless links on top of a wired NoC architecture. A variety of antennas and transceivers [29, 30, 31, 32, 33, 34, 35] have been proposed to be co-integrated with the processors, by considering the wireless channel inside the chip package as the propagation medium. WNoC complements NoC nicely with reconfigurability, low latency, and broadcasting capabilities for a high number of cores. In a dense wired routing network, WNoC can avoid unnecessary hops when communicating distant cores.
To meet the expected performance of WNoC with high data rates and ultra-low latency, the wireless channels have to be shared among the antennas or the modulation order data transmissions have to be increased. Therefore MAC protocols play an important role in maintaining reliable data communication, where several attempts have been made [36, 17, 10, 37, 38] to adapt token-passing, CSMA, time division multiplexing, code division multiple access to the chip environment by considering its unique requirements and characteristics. On the other hand, although the modulations are also crucial on enhancing the data rate, the area, power overhead of the chip scale transceivers and stringent BER requirements have limited the modulations to On-Off keying (OOK) [39, 6].
@@ II-B Wireless Channel within the Chip Package
The controlled materials and dimensions within the chip package let us tailor the wireless channel and know the nature of the medium beforehand [40, 13, 41]. A computer chip is created by stacking layers of lossy silicon, metal, and dielectric [32]. In particular, the structure of the flip-chip configurations widely used in nowadays processors is, from top to bottom: a heat sink on top, the silicon substrate, a thin dielectric layer for the interconnections, and a layer of solder bumps that connect the chip to the PCB [42]. Recently, a trend in computer architecture consists in the interconnection of multiple flipped chips in a co-planar configuration. To this end, the chips are connected through a silicon interposer with a much higher connection pin density than via a PCB [43, 44]. All the work developed in this paper considers such a 2.5D integration scheme.
Regardless of the number of interconnected chips, communicating within the chip package with electromagnetic radiation has its unique characteristics. A large amount of the transmitted energy will be reflected by the metal enclosure, the edges, and obstacles in the structure, leading to a large delay spread with multi-path propagation while creating a distinctive CIR between each transmitter and receiver [40, 13]. Fortunately, the structure is designed beforehand, and all materials, nodes, and terminals will become fixed and known. This offers a significant advantage since the wireless channel can be treated as quasi-deterministic and time-invariant which enables a pre-characterization of the CIR that will not change and that requires a single estimation process and can be re-used moving forward.
@@ II-C Time Reversal for On-Chip Communications
TR is a technique that can be used to compensate for the multipath effects that occur in complex transmission mediums with the use of adaptive energy focusing on the pre-characterized CIRs. The phase conjugate of the CIR is obtained, pre-emphasized with the transmitting symbols, and propagated through the same wireless medium such that the radiated energy is focused in both time and space at a target receiver. TR acts as an ideal spatial matched filter with the process of auto-correlation of the transmitted time-reversed channel response with an equal wireless medium. As the chip-scale wireless communications are conducted in quasi-deterministic channels, TR is a promising technology to adapt as a solution to mitigate the ISI effects that occur due to energy reverberation. With prior known wireless channels, repetitive CIR characterization is not required.
Channel correlation plays a crucial role in TR processing. Even though TR was initially adapted for on-chip communications to mitigate ISI, later it was proved that TR could be used to reduce the CCI in same frequency/time signal transmissions [27, 28]. The dispersive nature of the wireless medium inside the chip create unique wireless channels for each link. When these communicating channels are less correlated TR is capable of reducing the received energy to the nearby nodes, with the destructive interference occur due to high cross-correlation in between the channels. In return, it enables the wireless medium to have parallel concurrent signal transmissions by sharing the same bandwidth, thus improving the aggregate data rate of the system. However, this phenomenon is adverse with symmetric channels where the CCI could be high. In such cases, the known CIR can be analyzed before the transmissions to optimize the communication scheme. Nevertheless, the symmetric channels can be highly useful to focus the energy on multiple nodes with single transmitter excitation to enhance the broadcasting capabilities.
@@ III-A TR with a Single Receiver
where denotes the autocorrelation function of the channel and is the Additive White Gaussian Noise (AWGN) component at the receiver . The autocorrelation property eliminates the multipath effects with constructive and destructive interference cancellation while improving the signal-to-noise ratio (SNR) at the respective receiver.
@@
With the above expression, it is observed that when the transport signal is passed through symmetric channels (autocorrelation index is high) the energy components are constructively added with a null phase shift while having the maximum energy concentration at the point of maximum propagation delay; . When the TR adopted CIR is dissimilar, the cross-correlation index is phase shifted and the energy concentration will be controlled by the shifted amount of phase. As the waves are propagated through an enclosed chamber, it is complex to predict the randomness of the CIR. Thus, it is important to measure each CIR before the transmissions and design the protocols to optimize the paths in parallel communications.
@@ IV-A Full-wave Simulation
Finally, we note that, to validate the time reversal approach, we record the CIR and excite the transmitting antenna with the reversed version of the CIR. This simulation is used to obtain the field distribution and see the spatiotemporal focusing, as well as to later compare the results with that of a simple convolution between the filter and the actual channel in MATLAB.
@@ IV-B PHY Layer Simulation
At the receiver, energy detection is implemented by integrating the received signals over the respective time intervals based on the bandwidth. The computed energy values are compared with a threshold to decode the data stream. The threshold value is decided a posteriori and on a per-link basis to minimize the Bit Error Rate (BER).
@@ IV-C SINR Analysis with TR
To quantify the impact of TR on the effectiveness of the signal transmission, the signal-to-noise plus interference ratio (SINR) shall be assessed for both single-link and multi-link TR precoded communications.
We detect the peak energy concentration inside a defined time window with general energy detection [46]. The length of the time window () is obtained based on the expected bandwidth of the signal transmission. The energy is measured by placing the time window such that the maximum energy concentration occurs at and the energy components outside the defined time duration is considered as interference due to multi-path effects.
The SINR is defined for two cases as aforementioned in Section III-A and Section III-B. Therefore, SINR for single transmitter single receiver channel can be obtained as,
where , , , .
@@
Where , , .
@@ V-A Impact of Single-link Transmissions
With respect to CCI, an important conclusion can be drawn from Fig. 5(c). This image shows the ability of TR to compress the energy concentration in non-desired targets if and when the channel impulse response of each wireless path is not correlated with the targeted receiver. In this particular case, node is placed in a relatively symmetric position with respect to the targeted node and, hence, receives a relatively large interference. An opposite behavior is observed for nodes to , for which the interference is largely suppressed.
In Fig. 5(d) we compare the BER on TR and non-TR single-link transmissions. It is observed that with the superior spatialtemporal focusing in high dates TR shows a better link quality by reducing the ISI. However in non-TR transmissions, the BER is increased when the data rate is reaching to a few Gb/s with the high delay spread of the channel as observed in Fig. 7(a).
@@ V-B Impact of Multi-link Transmissions
To evaluate TR with parallel transmissions, first we ascess the near field coupling effects of the closely placed monopole antennas as shown in Fig. 3. In terms of mutual coupling, even though the antennas are located in a close proximity to each other, we observe in Fig. 6(a) the spatiotemporal focusing exhibit with TR in simultaneous transmission is similar with sequential individual TR transmissions. In here, we measure the ratio of average received energy at nodes and , with respect to the single link TR transmissions and multi-link signal transmissions from the nodes and . Thus, we observe that with the energy ratios the spatial isolation offer with TR precoded signals enable us to transmit multiple signals from multiple receivers at the same time, despite of the near-field impairments.
The energy focusing with TR precoded multi-link transmissions were measured with respect to received signal strength with a total of 10dBm transmit power in Fig. 6(b). The SINR of a single-transmitter-single-receiver and multiple-transmitter-multiple-receiver cases, as mentioned in Equation (10), Equation (11) were obtained respectively. It was observed that, when the transmission power is swept, an 8 dB SINR gap in between TR and non-TR signal transmissions appears. Based on the positions and the distance between the links CD and EF, there is approximately 1 dB gap in the SINR. However, it is illustrated that, in all cases, transmissions with TR have high SINR with focused signal and reduced multi-path effects. In summary, these results are indicative of the capacities of TR.
Furthermore, the performance of TR with multiple transmissions were measured in terms of BER as a function of the aggregated data rate (i.e. the sum of the data rates of all parallel transmissions), as illustrated in Fig. 6(c). It was observed that even with three concurrent communications, the focused energy is higher, and thus the BER is improved at the receiver. In terms of data rate, a significantly high aggregate data rate can be obtained with time reversal. In particular, a speed of 50 Gb/s is obtained with an error rate below 0.01 with a single-link transmission. Compared with the non-TR transmission, the TR case achieves an order of magnitude higher speed for the same BER. Further, as we add new concurrent links to the simulation, we observe that the aggregate data rate increases with the number of deployed, although the system starts to saturate with three concurrent links in this particular case.

### v3

@@ Abstract
The concept of Wireless Network-on-Chip (WNoC) has emerged as a potential solution to address the escalating communication demands of modern computing systems due to its low-latency, versatility, and reconfigurability. However, for WNoC to fulfill its potential, it is essential to establish multiple high-speed wireless links across chips. Unfortunately, the compact and enclosed nature of computing packages introduces significant challenges in the form of Co-Channel Interference (CCI) and Inter-Symbol Interference (ISI), which not only hinder the deployment of multiple spatial channels, but also severely restrict the symbol rate of each individual channel. In this paper, we posit that Time Reversal (TR) could be effective in addressing both impairments in this static scenario, thanks to its spatiotemporal focusing capabilities even in the near-field. Through comprehensive full-wave simulations and bit error rate analysis in multiple chip layouts with multiple frequency bands, we provide evidence that TR can increase the symbol rate by an order of magnitude, enabling the deployment of multiple concurrent links and achieving aggregate speeds exceeding 100 Gb/s. Finally, we evaluate the impact of reducing the sampling rate of the TR filter on the achievable speeds, paving the way to practical TR-based wireless communications at the chip scale.
I Introduction
@@
In this context, Wireless Network-on-Chip (WNoC) is an emerging paradigm with significant promise to enhance communications within computing packages. WNoC are possible because of the advances in antennas and transceivers in the millimeter-wave (mmWave) and terahertz (THz) bands that have made possible their integration inside computing cores [3]. They are expected to mitigate the scalability and flexibility issues that come with the increment of the number of processors, their integration on multiple interconnected chips, and the use of wired NoC communication approaches [1, 4]. The goal is to augment the wired backbone with wireless links within the package [5].
@@
Due to the use of electromagnetic (EM) propagation nearly at the speed of light, WNoC provides broadcast and multicast opportunities with reduced latency and is expected to offer reconfigurable links able to operate at speeds of tens of Gb/s or more. For the WNoC paradigm to be competitive, though, it must deliver a reliability commensurate to that of their wired counterpart. In particular, multiple works [6, 7, 8] point towards a latency of 1–10 nanoseconds, a throughput in the range of 10–100 Gb/s or higher, and a Bit Error Rate (BER) of 10-12–10-15 as potential targets for WNoC. Complying with these values while keeping low the area/power consumption calls for affordable solutions, low order modulations, and simple physical layer elements for implementation [9, 7, 10].
@@
However, several challenges arise when including wireless communications in a package environment while trying to meet the target performance requirements. Some examples include the antenna placement in the highly integrated environment, design of transceivers with efficient coding schemes and modulations, interference management, and medium access control [5, 11, 12]. In particular, in this paper we focus on the problems of Inter-Symbol Interference (ISI) and Co-Channel Interference (CCI). In the on-chip scenario, the antennas radiate in a highly reverberant environment and are impaired by near-field effects. As a result, on the one hand, signals are subject to multiple reflections before reaching the receiver, leading to large delay spreads dominated by the strong multipath components. This leads to severe ISI, which degrades the performance of high-rate transmissions especially in low-order modulations. On the other hand, reverberation causes the radiated energy to spread across the entire space. This leads to strong CCI with non-intended antennas and prevents the possibility of having multiple spatial channels concurrently. Other techniques such as Orthogonal Frequency Division Multiple Access (OFDMA) and beamforming could address ISI and CCI, but are not advisable in the chip-scale communications scenario due to the required complex and power-hungry transceivers, as well as the need for multiple antennas consuming a large area, respectively [13, 14].
@@
Time Reversal (TR) [15, 16, 17, 18] is a technique that uses a time-reversed version of the channel’s impulse response (CIR) to create a matched filter that focuses energy in space and time in a targeted destination. Due to the spatiotemporal focusing, TR can therefore considerably alleviate both ISI and CCI, as shown in multiple works at multiple scales and including in the near-field [19, 16].
@@
The rest of the paper is organized as follows. In Section II, we provide some background related to the general problems of the on-chip scenario. In Section III, we formulate the system model and in Section IV, we provide a detailed description of the methodology, including the simulation pipeline and the evaluated variations. We present the results obtained for the parametric characterization from the base scenario in Section V, while Section VI evaluates the performance of TR with the changes in the antenna operating frequency and the geometry of the chip. Finally, in Section VII, we assess the trade-off between performance and feasibility of TR filters and Section VIII discuss the limitations. Section IX summarize the contribution and outlines directions for future work.
@@ II-A Wireless Communications at the Chip Scale
Wireless on-chip communications have been built with the foundation of using mmwave technology, as an overlay of wireless links on top of a wired NoC architecture. A variety of antennas and transceivers [22, 23] have been proposed to be co-integrated with the processors, by considering the wireless channel inside the chip package as the propagation medium. WNoC complements NoC nicely with reconfigurability, low latency, and broadcasting capabilities for a high number of cores. In a dense wired routing network, WNoC can avoid unnecessary hops when communicating with distant cores.
To meet the expected performance of WNoC with high data rates and ultra-low latency, the wireless channels have to be shared among the antennas or the modulation order of data transmissions have to be increased. Therefore MAC protocols play an important role in maintaining reliable data communication, where several attempts have been made [24, 25, 8] to adapt token-passing, CSMA to the chip environment by considering its unique requirements and characteristics. On the other hand, although the modulations are also crucial on enhancing the data rate, the area, power overhead of the chip scale transceivers and stringent BER requirements have limited the modulations to On-Off keying (OOK) [26, 5].
@@ II-B Wireless Channel within the Chip Package
The controlled materials and dimensions within the chip package let us tailor the wireless channel and know the nature of the medium beforehand [10, 27]. A computer chip is created by stacking layers of lossy silicon, metal, and dielectric [28]. In particular, the structure of the flip-chip configurations widely used in nowadays processors is, from top to bottom: a heat sink on top, the silicon substrate, a thin dielectric layer for the interconnections, and a layer of solder bumps that connect the chip to the PCB [29]. Recently, a trend in computer architecture consists in the interconnection of multiple flipped chips in a co-planar configuration. To this end, the chips are connected through a silicon interposer with a much higher connection pin density than via a PCB [30, 31]. All the work developed in this paper considers such a 2.5D integration scheme.
Regardless of the number of interconnected chips, communicating within the chip package with EM radiation has its unique characteristics. A large amount of the transmitted energy will be reflected by the metal enclosure, the edges, and obstacles in the structure, leading to a large delay spread with multipath propagation while creating a distinctive CIR between each transmitter and receiver [10]. Fortunately, the structure is designed beforehand, and all materials, nodes, and terminals will become fixed and known. This offers a significant advantage since the wireless channel can be treated as quasi-deterministic and time-invariant which enables a pre-characterization of the CIR that will not change and that requires a single estimation process and can be re-used moving forward.
@@ II-C Time Reversal for On-Chip Communications
TR is a technique that can be used to compensate for the multipath effects that occur in complex transmission mediums with the use of adaptive energy focusing on the pre-characterized CIRs. The phase conjugate of the CIR is obtained, pre-emphasized with the transmitting symbols, and propagated through the same wireless medium such that the radiated energy is focused in both time and space at a target receiver as shown in Fig. 2. TR acts as an ideal spatial matched filter with the process of auto-correlation of the transmitted time-reversed channel response with an equal wireless medium. As the chip-scale wireless communications are conducted in quasi-deterministic channels, TR is a promising technology to adapt as a solution to mitigate the ISI effects that occur due to energy reverberation. With prior known wireless channels, repetitive CIR characterization is not required.
Channel correlation plays a crucial role in TR processing. Even though TR was initially adapted for on-chip communications to mitigate ISI, later it was proved that TR could be used to reduce the CCI in same frequency/time signal transmissions [20, 21]. The dispersive nature of the wireless medium inside the chip create unique wireless channels for each link. When these communicating channels are less correlated TR is capable of reducing the received energy to the nearby nodes, with the destructive interference occur due to low cross-correlation in between the channels. In return, it enables the wireless medium to have parallel concurrent signal transmissions by sharing the same bandwidth, thus improving the aggregate data rate of the system. However, this phenomenon is adverse with symmetric channels where the CCI could be high. In such cases, the known CIR can be analyzed before the transmissions to optimize the communication scheme. Nevertheless, the symmetric channels can be highly useful to focus the energy on multiple nodes with single transmitter excitation to enhance the broadcasting capabilities.
@@ III-A TR with a Single Receiver
where denotes the autocorrelation function of the channel and is the Additive White Gaussian Noise (AWGN) component at the receiver . With an ideal TR filter response, the autocorrelation property eliminates the multipath effects with constructive and destructive interference cancellation and gains strong temporal focusing while improving the signal-to-interference plus noise ratio (SINR) at the respective receiver. Yet, the cross-correlation between the transmitter-receiver channel and the channels at nearby locations will affect the degree of spatial focusing and residual side lobes may persist.
@@
With the above expression, it is observed that when the transport signal is passed through symmetric channels (auto correlation index is high) the energy components are constructively added with a null phase shift while having the maximum energy concentration at the point of maximum propagation delay; . When the TR adopted for dissimilar CIR, the cross-correlation index is phase shifted and the energy concentration will be controlled by the shifted amount of phase.
@@ IV-A Full-wave Simulation
Finally, we note that, to validate the TR approach, we record the CIR and excite the transmitting antenna with the reversed version of the CIR. This simulation is used to obtain the field distribution and see the spatiotemporal focusing, as well as to later compare the results with that of a simple convolution between the filter and the actual channel in MATLAB.
@@ IV-B PHY Layer Simulation
At the receiver, energy detection is implemented by integrating the received signals over the respective time intervals based on the bandwidth. The computed energy values are compared with a threshold to decode the data stream. The threshold value is decided a posteriori and on a per-link basis to minimize the BER. The BER values represent pre-Forward Error Correction (FEC) performance of the links.
@@ IV-C SINR Analysis with TR
To quantify the impact of TR on the effectiveness of the signal transmission, the SINR shall be assessed for both single-link and multi-link TR precoded communications.
We detect the peak energy concentration inside a defined time window with general energy detection [33]. The length of the time window () is obtained based on the expected bandwidth of the signal transmission. The energy is measured by placing the time window such that the maximum energy concentration occurs at and the energy components outside the defined time duration is considered as interference due to multipath effects.
The SINR is defined for two cases as aforementioned in Section III-A and Section III-B. Therefore, SINR for single transmitter single receiver channel can be obtained as
where , , , . Here, represents the signal energy of the received signal, while denotes the ISI caused due to multi path effects of the channel.
@@
where , , . denotes the energy of the received signal, captures the ISI due to multi path propagation and is the CCI due to multiple simultaneous transmissions.
@@ V-A Impact of Single-link Transmissions
With respect to CCI, an important conclusion can be drawn from Fig. 6(c). This image shows the ability of TR to compress the energy concentration in non-desired targets if and when the CIR of each wireless path is not correlated with the targeted receiver. In this particular case, node is placed in a relatively symmetric position with respect to the targeted node and, hence, receives a relatively large interference. An opposite behavior is observed for nodes to , for which the interference is largely suppressed.
In Fig. 6(d) we compare the BER on TR and non-TR single-link transmissions. Theoretical BER with Q-function for OOK modulation was compared with simulations. It is observed that with the superior spatial-temporal focusing in high data rates TR shows a better link quality by reducing the ISI. However in non-TR transmissions, the BER is increased when the data rate is reaching to a few Gb/s with the high delay spread of the channel as observed in Fig. 8(a).
@@ V-B Impact of Multi-link Transmissions
To evaluate TR with parallel transmissions, first we assess the near-field coupling effects of the closely placed monopole antennas as shown in Fig. 4. In terms of mutual coupling, even though the antennas are located in a close proximity to each other, we observe in Fig. 7(a) the spatiotemporal focusing shown with TR in simultaneous transmission is similar with sequential individual TR transmissions. In here, we measure the ratio of average received energy at nodes and , with respect to the single link TR transmissions and multi-link signal transmissions from the nodes and . Thus, we observe that with the energy ratios the spatial isolation offer with TR precoded signals enable us to transmit multiple signals from multiple receivers at the same time, despite of the near-field impairments.
The energy focusing with TR precoded multi-link transmissions were measured with respect to received signal strength with a total of 10 dBm transmit power in Fig. 7(b). The SINR of a single-transmitter-single-receiver and multiple-transmitter-multiple-receiver cases, as mentioned in Equation (10), Equation (11) were obtained respectively. It was observed that, when the transmission power is swept, an 8 dB SINR gap in between TR and non-TR signal transmissions appears. Based on the positions and the distance between the links CD and EF, there is approximately 1 dB gap in the SINR. However, it is illustrated that, in all cases, transmissions with TR have high SINR with focused signal and reduced multipath effects. In summary, these results are indicative of the capacities of TR.
Furthermore, the performance of TR with multiple transmissions were measured in terms of BER as a function of the aggregated data rate (i.e. the sum of the data rates of all parallel transmissions), as illustrated in Fig. 7(c). It was observed that even with three concurrent communications, the focused energy is higher, and thus the BER is improved at the receiver. In terms of data rate, a significantly high aggregate data rate can be obtained with time reversal. In particular, a speed of 50 Gb/s is obtained with an error rate below 0.001 with a single-link transmission. Compared with the non-TR transmission, the TR case achieves an order of magnitude higher speed for the same BER. Further, as we add new concurrent links to the simulation, we observe that the aggregate data rate increases with the number of deployed, although the system starts to saturate with three concurrent links in this particular case.